is f(x)=t^2+2t-3 an odd or even function? how?
Answer by stanbon(75887)

f(t)=t^2+2t-3 an odd or even function? how?
---------------
f(t)=t^2+2t-3
f(-t) = t^2-2t-3
-f(-t) = -t^2+2t+3
----
Since f(t) is not equal to f(-t), f is not even.
----
Since f(t) is not equal to -f(-t), f is not odd.
=============
f is neither even nor odd.
================
Cheers,
stan H.